Consider a computer system with a 32-bit logical address and 8-KB page size. The system supports up to 1GB of physical memory. How many entries are there in a page table?
Answer #1

Given computer system has 32-bit logical address

Logical Address Space size = 2^32 bytes

Physical Memory size = 1GB = 2^30 bytes

Page size = 8 KB = 2^3 * 2^10 bytes = 2^13 bytes

a)

Number of entries in single-level page table = Number of pages in the Logical Memory
  
= 2^32 bytes / 2^13 bytes = 2^19 entries

b)

Number of entries in inverted page table = number of pages/frames in the Physical memory

= 2^30 / 2^13 = 2^17 entries

PLEASE UPVOTE IF THE ANSWER WAS HELPFUL!!